DEMAND   April 71          May 66          June 91          July 71         &nbs

DEMAND
  April 71       
  May 66       
  June 91       
  July 71       
  August 96       
  September 91       

1. Using single exponential smoothing with α = 0.20 and a September forecast = 49, calculate a forecast for October.

2.Using simple linear regression, calculate the trend line for the historical data. Say the X axis is April = 1, May = 2, and so on, while the Y axis is demand. (Round your intercept value to the nearest whole number and slope value to 2 decimal places.)

3.
Calculate a forecast for October using your regression formula

PLEASE **STAR** YOUR FINAL ANSWER, THANK YOU
Answer #1

Answer 1=

Exponential smoothing forecasting = Ft = Ft-1 +α(At-1-Ft-1)

In the question, Ft-1 =49 At-1=91, α=0.2

Exponential smoothing forecasting for September = Ft = Ft-1 +α(At-1-Ft-1) =49+0.2*(91-49)=57.4

****Answer=57.4****

Answer 2=

x y xy x^2 y^2
1 71 71 1 5041
2 66 132 4 4356
3 91 273 9 8281
4 71 284 16 5041
5 96 480 25 9216
6 91 546 36 8281
Sum 21 486 1786 91 40216
Σν) Σr) - Σx ) Σxy) (Σχ2)-Σx > n (Σxν) - Σx ) (Σν ) η22)-Σx )> α
b= 85/(17.5*1450)^0.5)
b=0.5336
a=(1450/5)^0.5
b= (6*1789-21*486)/(6*91-21*21)
b=5.03
a= (486*91-21*1786)/(6*91-21*21)
a=64
Regression line y=64+5.03x

So the answer is

****y=64+5.03x*****

Answer 3= Forcast for October,

put x=7 in y=64+5.03x

y=64+5.03*7

y=99.21

Answer= ***99.21**